# DFALLM: Achieving Generalizable Multitask Deepfake Detection by Optimizing Audio LLM Components

### v1

@@ I Introduction
Audio deepfake detection has become a significant public concern regarding security and reliability. Traditional detection methods, often based on smaller audio models like WavLM [1], achieve high accuracy on known datasets. However, they suffer from two critical limitations. First, they lack generalizability; their performance degrades sharply when confronted with newly emerging spoofing techniques or out-of-domain (OOD) data [2]. Second, they are single-task models, struggling to handle multiple tasks like spoof attribution or localization within a single model.
@@
However, when directly applied to deepfake detection, existing ALLMs still exhibit a generalization bottleneck. We argue that this phenomenon is due to the information bottleneck in their audio encoders and corresponding textual LLMs [3].
Currently, most of ALLMs utilise Whisper [4] as their audio encoders [5, 6, 7]. However, Whisper is typically trained with supervision on ASR tasks. This training objective compels the model to preserve linguistic content while simultaneously discarding non-linguistic, acoustic details deemed irrelevant to speech recognition. Critically, it is precisely this discarded acoustic information that has been shown to be essential for robust deepfake detection.
This analysis leads to our core hypothesis: the key bottleneck for ALLM generalization is an encoder-level problem, not an LLM-level one. We therefore posit that the key to unlocking performance is not the LLM, but the choice of the encoder. To verify this, we conduct a systematic investigation of the core ALLM components, explicitly comparing semantic-optimized encoders (like Whisper) against acoustically-aware encoders (like Wav2Vec2-BERT [8]) which preserve more raw signal features through self-supervised masked modeling.
@@ II Related work
Current deepfake detection research can be grouped into three subtasks: binary detection, attribution, and localization. Binary detection focuses on distinguishing bona fide from spoofed audio, with early work relying on deep learning models such as Whisper or spectrogram-based classifiers [2]; AASIST [9] and post-training approaches [10] further advanced performance. Attribution identifies which generative model produced a deepfake [11], leveraging similar techniques, including digital fingerprint analysis [12] and large-scale benchmark studies [13]. Localization aims to determine the precise manipulated segments [14], with compact models proposed to capture fine-grained feature patterns [15]. Despite these developments, prior work typically addresses the three subtasks in isolation, remains largely unimodal, and continues to face challenges in generalization.
With the rise of ALLMs such as AudioPaLM [16], joint speech–text reasoning has shown strong results in ASR and audio captioning, yet their use in forensic audio, particularly deepfake detection, remains limited [3]. Prior work suggests that fine-tuning the audio encoder is key, rather than keeping it frozen, but evaluations have focused solely on ASVspoof 2019, with no OOD testing, leaving generalizability unclear.
ALLM has the potential of multitask learning and generalization with large number of parameters. Previous ALLMs have demonstrated strong performance across a wide range of audio understanding tasks, such as speech recognition, audio captioning, and spoken question answering [17, 18]. A key factor to their success is their integration of textual LLMs, such as Qwen 2 [19] and LLaMA [20], as foundational components. These LLMs bring knowledge and reasoning and generalization abilities. ALLMs can hence effectively bridge acoustic inputs with high-level language semantics. However, multitask learning in deepfake analysis that combines detection, attribution, and localization has been little explored.
Therefore, the primary aim of this paper is to investigate the capability of ALLMs in deepfake detection. In addition, we explore an important open question concerning the scaling behavior of large models [21], which suggests that model capacity and data size jointly determine overall performance in large language models. Our work empirically examines these factors to identify the optimal model scale, as their influence on the deepfake detection ability of ALLMs remains insufficiently understood.
@@ IV-A Datasets
Training Corpus: We built a combined training dataset by sampling from multiple public deepfake audio datasets, including ASVspoof 2019 [22], SpoofCeleb [23], MLAADv6 [24], ReplayDF [25], DFADD [26], AISHELL3 [27], ADD2023 [28], GigaSpeech [29], CNCeleb [30], PartialSpoof [31].
Evaluation Sets: Evaluation Sets: We used distinct evaluation sets for each task. For Detection: To assess both in-domain and out-of-domain (OOD) performance, we employed three datasets. ASVSpoof 2019 LA (71,237 samples) serves as the in-domain set . In-the-Wild (ITW) [32] (31,780 samples) and Demopage (2,182 aggregated samples) are used as OOD test sets. These are provided in supplementary materials, with Demopage meaning the collection from samples shown on the demo page of various TTS models.
@@ IV-C Training and Implementation Details
A projection module is used to map the audio representations from the audio encoder into the text embedding space. We use a Linear layer as the projection module. For our training strategy, we fully trained the audio encoder and projector to adapt it to the acoustic features of the task. The LLM was then efficiently fine-tuned using Low-Rank Adaptation (LoRA) [33].
@@
We experimented with multiple audio encoders, including Wav2Vec2-Bert [8] and Whisper [4], to extract robust acoustic representations from speech signals. For the textual component, we explored different LLMs, including Qwen2.5 [19], Qwen3 [34] and LLaMA [35], to process both the projected audio representations and the accompanying textual prompts.
We selected these models because the chosen audio encoders are representative of current architectures, each emphasizing different aspects of audio understanding. Whisper prioritizes semantic comprehension, as it is trained primarily on large-scale ASR corpora, whereas Wav2Vec2-BERT, derived from Wav2Vec2 [36], leverages self-supervised masked language modeling, thereby preserving more of the raw acoustic and physical characteristics of the speech signal. Comparing these two encoders allows us to identify potential bottlenecks in audio representation that may hinder the generalizability of ALLMs in deepfake detection. For the textual LLMs, we selected two models from the Qwen family and one from the LLaMA family to maintain diversity in the model pool while enabling controlled comparisons across architectures with differing linguistic priors and adaptation strategies.
@@ V-A Main findings
Although the aforementioned ALLM configurations demonstrate some generalizability on out-of-distribution datasets, it remains necessary to evaluate their performance across multiple tasks, including attribution and localization. Multitask evaluation provides a rigorous test of whether the models truly understand and analyze the audio signals that indicate a deepfake, rather than relying on spurious correlations or chance [2]. A more comprehensive assessment is therefore required to determine if the models can accurately identify both the presence of a deepfake and its underlying characteristics. Our objective is to empirically assess whether ALLMs achieve superior performance not only on individual tasks but also in a multitask setting encompassing attribution and localization.
@@ V-B Ablation study
First, audio representations are influenced not only by the choice of the audio encoder but also by the characteristics of the original audio input. Previous explainable AI studies have shown that certain distinguishable deepfake features are concentrated within specific frequency ranges of the audio [37, 38, 39]. Moreover, different audio encoders operate at varying frame rates, which affect the effective frequency resolution of the input; for example, Whisper uses a frame rate of 12.5 Hz, whereas Wav2Vec2 operates at 50 Hz. Consequently, we also investigate the optimal framerate as a component of our overall configuration, with results shown in Table IV.
@@
These results reflect a behavior that fits for LLM scaling law [40]: performance improves steadily as more training data is provided, even though the underlying LLM remains small. It also indicates that achieving robust performance with LLMs is not straightforward, as the task requires sufficient data for the model to fully understand the audio and accurately identify potential deepfaked segments. While smaller models can perform reasonably well on relatively constrained datasets, larger models are able to realize their full potential when provided with more extensive training data. This may be because deepfake detection demands a genuine understanding of the audio content rather than superficial signal-based guesses. These findings further suggest that ALLMs are capable of learning intrinsic audio features that support effective detection across varying conditions.

### v2

@@ I Introduction
Audio deepfake detection has become a significant public concern regarding security and reliability. Traditional detection methods, often based on smaller audio models like WavLM [Chen_2022], achieve high accuracy on known datasets. However, they suffer from two critical limitations. First, they lack generalizability; their performance degrades sharply when confronted with newly emerging spoofing techniques or out-of-domain (OOD) data [yi2023audio]. Second, they are single-task models, struggling to handle multiple tasks like spoof attribution or localization within a single model.
@@
However, when directly applied to deepfake detection, existing ALLMs still exhibit a generalization bottleneck. We argue that this phenomenon is due to the information bottleneck in their audio encoders and corresponding textual LLMs [gu2025allm4addunlockingcapabilitiesaudio].
Currently, most of ALLMs utilise Whisper [radford2022robustspeechrecognitionlargescale] as their audio encoders [chu2023qwenaudioadvancinguniversalaudio, ding2025kimi, alex2025palprobingaudioencoders]. However, Whisper is typically trained with supervision on ASR tasks. This training objective compels the model to preserve linguistic content while simultaneously discarding non-linguistic, acoustic details deemed irrelevant to speech recognition. Critically, it is precisely this discarded acoustic information that has been shown to be essential for robust deepfake detection.
This analysis leads to our core hypothesis: the key bottleneck for ALLM generalization is an encoder-level problem, not an LLM-level one. We therefore posit that the key to unlocking performance is not the LLM, but the choice of the encoder. To verify this, we conduct a systematic investigation of the core ALLM components, explicitly comparing semantic-optimized encoders (like Whisper) against acoustically-aware encoders (like Wav2Vec2-BERT [chung2021w2vbertcombiningcontrastivelearning]) which preserve more raw signal features through self-supervised masked modeling.
@@ II Related work
Current deepfake detection research can be grouped into three subtasks: binary detection, attribution, and localization. Binary detection focuses on distinguishing bona fide from spoofed audio, with early work relying on deep learning models such as Whisper or spectrogram-based classifiers [yi2023audio]; AASIST [Jung2021AASIST] and post-training approaches [ge2025posttrainingdeepfakespeechdetection] further advanced performance. Attribution identifies which generative model produced a deepfake [Klein_2024], leveraging similar techniques, including digital fingerprint analysis [Yan_2022] and large-scale benchmark studies [xie2025neuralcodecsourcetracing]. Localization aims to determine the precise manipulated segments [he2025manipulated], with compact models proposed to capture fine-grained feature patterns [zhang2024mfms]. Despite these developments, prior work typically addresses the three subtasks in isolation, remains largely unimodal, and continues to face challenges in generalization.
With the rise of ALLMs such as AudioPaLM [rubenstein2023audiopalmlargelanguagemodel], joint speech–text reasoning has shown strong results in ASR and audio captioning, yet their use in forensic audio, particularly deepfake detection, remains limited [gu2025allm4addunlockingcapabilitiesaudio]. Prior work suggests that fine-tuning the audio encoder is key, rather than keeping it frozen, but evaluations have focused solely on ASVspoof 2019, with no OOD testing, leaving generalizability unclear.
ALLM has the potential of multitask learning and generalization with large number of parameters. Previous ALLMs have demonstrated strong performance across a wide range of audio understanding tasks, such as speech recognition, audio captioning, and spoken question answering [das2024speechverse, yang2024uniaudio]. A key factor to their success is their integration of textual LLMs, such as Qwen 2 [team2024qwen2] and LLaMA [touvron2023llama], as foundational components. These LLMs bring knowledge and reasoning and generalization abilities. ALLMs can hence effectively bridge acoustic inputs with high-level language semantics. However, multitask learning in deepfake analysis that combines detection, attribution, and localization has been little explored.
Therefore, the primary aim of this paper is to investigate the capability of ALLMs in deepfake detection. In addition, we explore an important open question concerning the scaling behavior of large models [held2025relative], which suggests that model capacity and data size jointly determine overall performance in large language models. Our work empirically examines these factors to identify the optimal model scale, as their influence on the deepfake detection ability of ALLMs remains insufficiently understood.
@@ IV-A Datasets
Training Corpus: We built a combined training dataset by sampling from multiple public deepfake audio datasets, including ASVspoof 2019 [wang2020asvspoof], SpoofCeleb [jung2025spoofcelebspeechdeepfakedetection], MLAADv6 [müller2025mlaadmultilanguageaudioantispoofing], ReplayDF [muller2025replaydf], DFADD [du2024dfadddiffusionflowmatchingbased], AISHELL3 [shi2021aishell3multispeakermandarintts], ADD2023 [yi2023add2023secondaudio], GigaSpeech [Chen_2021], CNCeleb [fan2019cncelebchallengingchinesespeaker], PartialSpoof [Zhang_2023].
Evaluation Sets: Evaluation Sets: We used distinct evaluation sets for each task. For Detection: To assess both in-domain and out-of-domain (OOD) performance, we employed three datasets. ASVSpoof 2019 LA (71,237 samples) serves as the in-domain set . In-the-Wild (ITW) [muller2022does] (31,780 samples) and Demopage (2,182 aggregated samples) are used as OOD test sets. These are provided in supplementary materials, with Demopage meaning the collection from samples shown on the demo page of various TTS models.
@@ IV-C Training and Implementation Details
A projection module is used to map the audio representations from the audio encoder into the text embedding space. We use a Linear layer as the projection module. For our training strategy, we fully trained the audio encoder and projector to adapt it to the acoustic features of the task. The LLM was then efficiently fine-tuned using Low-Rank Adaptation (LoRA) [hu2021loralowrankadaptationlarge].
@@
We experimented with multiple audio encoders, including Wav2Vec2-Bert [chung2021w2vbertcombiningcontrastivelearning] and Whisper [radford2022robustspeechrecognitionlargescale], to extract robust acoustic representations from speech signals. For the textual component, we explored different LLMs, including Qwen2.5 [team2024qwen2], Qwen3 [yang2025qwen3technicalreport] and LLaMA [touvron2023llamaopenefficientfoundation], to process both the projected audio representations and the accompanying textual prompts.
We selected these models because the chosen audio encoders are representative of current architectures, each emphasizing different aspects of audio understanding. Whisper prioritizes semantic comprehension, as it is trained primarily on large-scale ASR corpora, whereas Wav2Vec2-BERT, derived from Wav2Vec2 [baevski2020wav2vec20frameworkselfsupervised], leverages self-supervised masked language modeling, thereby preserving more of the raw acoustic and physical characteristics of the speech signal. Comparing these two encoders allows us to identify potential bottlenecks in audio representation that may hinder the generalizability of ALLMs in deepfake detection. For the textual LLMs, we selected two models from the Qwen family and one from the LLaMA family to maintain diversity in the model pool while enabling controlled comparisons across architectures with differing linguistic priors and adaptation strategies.
@@ V-A Main findings
Although the aforementioned ALLM configurations demonstrate some generalizability on out-of-distribution datasets, it remains necessary to evaluate their performance across multiple tasks, including attribution and localization. Multitask evaluation provides a rigorous test of whether the models truly understand and analyze the audio signals that indicate a deepfake, rather than relying on spurious correlations or chance [yi2023audio]. A more comprehensive assessment is therefore required to determine if the models can accurately identify both the presence of a deepfake and its underlying characteristics. Our objective is to empirically assess whether ALLMs achieve superior performance not only on individual tasks but also in a multitask setting encompassing attribution and localization.
@@ V-B Ablation study
First, audio representations are influenced not only by the choice of the audio encoder but also by the characteristics of the original audio input. Previous explainable AI studies have shown that certain distinguishable deepfake features are concentrated within specific frequency ranges of the audio [maltby2024frequency, feng2024audios, li2024detectingmachinegeneratedmusicexplainability]. Moreover, different audio encoders operate at varying frame rates, which affect the effective frequency resolution of the input; for example, Whisper uses a frame rate of 12.5 Hz, whereas Wav2Vec2 operates at 50 Hz. Consequently, we also investigate the optimal framerate as a component of our overall configuration, with results shown in Table IV.
@@
These results reflect a behavior that fits for LLM scaling law [kaplan2020scalinglawsneurallanguage]: performance improves steadily as more training data is provided, even though the underlying LLM remains small. It also indicates that achieving robust performance with LLMs is not straightforward, as the task requires sufficient data for the model to fully understand the audio and accurately identify potential deepfaked segments. While smaller models can perform reasonably well on relatively constrained datasets, larger models are able to realize their full potential when provided with more extensive training data. This may be because deepfake detection demands a genuine understanding of the audio content rather than superficial signal-based guesses. These findings further suggest that ALLMs are capable of learning intrinsic audio features that support effective detection across varying conditions.